What are the next two terms in this sequence?
15, 30, 20, 27, 25, 24, 30, 21, 35, 18, ___, ___

Answer by checkley77(12569)

This sequence requires a 2 level solution.
15, 30, 20, 27, 25, 24, 30, 21, 35, 18, 6, 15 First find the difference.
+15, -10, +7, -2, -1, +6, -9, -14,-17, Now find another level of differences.
????+5, -3, +5, -3, +5, -3, +5, -3, +5, -3,
Calculate the first missing number:
-17+5=-12+18=6 ans.
Calculate the next number:
18-6=12-3=9+6=15 ans.